# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ B.2.3 Technical Lemmas
Lemma B.1.
Let $(a_{k})_{k\in\mathbb{N}}\subset\mathbb{R}_{+}$ be a sequence of nonnegative real numbers with $a_{0}\in\mathbb{R}_{++}$. Then, the following inequalities hold:
@@
We state without proof a similar descent lemma for $(L_{0},L_{1})$-smooth functions. Its proof can be found in Zhang et al. (2020a, Lemma A.3).
Lemma B.2 (Descent lemma for $(L_{0},L_{1})$-smooth functions).
@@
B.3 Convergence Results for AdaGrad-Norm and AdaGrad
B.3.1 Proof of Theorem 5.1
@@ B.3 Convergence Results for AdaGrad-Norm and AdaGrad
$=\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle,$ (B.4)
@@
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (B.5)
@@
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle\\ \leqslant\frac{1}{2}\left(1+\frac{1}{\rho}\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{\tau L^{2}\alpha^{2}}{2}(1+\rho\tau)\frac{\left\lVert g_{k}\right\rVert^{2}}{v_{k+1}^{\nicefrac{{3}}{{2}}}}.$ (B.6)
Proof of Lemma B.3.
@@
$\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}=\frac{v_{k}-v_{k-1}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}=\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}.$ (B.7)
@@
$\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]$ $\leqslant\mathbb{E}_{k}\left[\frac{\left\lVert\nabla F(x_{k-1})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k}}}\right]$
$\qquad+\frac{1}{\sqrt{v_{k-1}}}\left(L^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2L\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\right)$
@@
By the A.M.-G.M. inequality with $\rho\in\mathbb{R}_{++}$, we have
@@
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle$ $\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{2\sqrt{v_{k-1}}}+\frac{\tau}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{\rho\tau^{2}L^{2}\alpha^{2}}{2}\cdot\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2\rho}\cdot\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau L^{2}\alpha^{2}}{2}\cdot\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}$
@@
We now continue with the proof of Theorem 5.1. Using Lemma B.3, (B.5) is upper bounded by
@@
$\mathbb{E}[F(x_{k+1})]\leqslant\mathbb{E}[F(x_{k})]-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\alpha}{2}\left(1-\frac{1}{\rho}\right)\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]+\frac{\tau\alpha}{2}\mathbb{E}[\varphi_{k-1}-\varphi_{k}]\\ +\frac{L\alpha^{2}}{2}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right].$ (B.8)
Summing the last two terms on the right hand side over $k$ from $1$ to $K$ and using Lemma B.1 yields
$\frac{L\alpha^{2}}{2}\sum_{k=1}^{K}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right]\leqslant\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}}.$
Thus, summing (B.8) over $k$ from $1$ to $K$ and rearranging yields
@@
Lemma B.4.
For any positive constants $(a,b)\in\mathbb{R}_{++}^{2}$, if $x\leqslant a+b\log x$ for any $x\in\mathbb{R}_{++}$, then $x\leqslant 2a-2b+4b\log(b/2+1)\leqslant 2a+4b\log(b/2+1)$.
Proof of Lemma B.4.
Let $h(x)\coloneqq x/2-\log x$. Then $h^{\prime}(x)=1/2-b/x=0$ if and only if $x=2b$. Since $h^{\prime\prime}(x)=b/x^{2}>0$ for any $x\in\mathbb{R}_{++}$, so $h$ achieves its minimum at $2b$, which means $h(x)\geqslant h(2b)=b-b\log(2b)$. By the A.M.-G.M. inequality, we have $b+2\geqslant 2\sqrt{2b}$ and taking logarithm yields $\log(b+2)\geqslant\log 2+\frac{1}{2}(\log 2+\log b)$. Therefore, we obtain $b\log(2b)=b(\log 2+\log b)\leqslant 2b(\log(b+2)-\log 2)=2b\log\left(b/2+1\right)$. Consequently, we get $x-b\log x=x/2+x/2-b\log x\geqslant x/2+b-b\log(2b)\geqslant x/2+b-2b\log\left(b/2+1\right)$. Therefore, we have $x/2+b-2b\log\left(b/2+1\right)\leqslant a$, which implies the desired inequality. ∎
By Lemma B.4, we have $\mathbb{E}[\sqrt{v_{k}}]\leqslant 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)\eqqcolon c_{3}$.
@@
$c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{k}}]\geqslant\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\geqslant\mathbb{E}\left[\frac{1}{\sqrt{v_{K}}}\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}\right]\geqslant\frac{\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{2}}{\mathbb{E}[\sqrt{v_{K}}]},$
@@
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\frac{1}{\sqrt{v_{k-1}}}\odot\nabla F(x_{k})\right\rangle\\ +\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)\odot g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{1}{\sqrt{v_{k}}}\odot g_{k}\right\rVert^{2}.$ (B.9)
@@
$\mathbb{E}_{k}\left[\partial_{j}F(x_{k})\left(\frac{1}{\sqrt{v_{k-1,j}}}-\frac{1}{\sqrt{v_{k,j}}}\right)g_{k,j}\right]\leqslant\frac{1}{2}\left(1+\frac{1}{\rho}\right)\frac{(\partial_{j}F(x_{k}))^{2}}{\sqrt{v_{k-1,j}}}+\frac{\tau}{2}\mathbb{E}_{k}[\widetilde{\varphi}_{k-1,j}-\widetilde{\varphi}_{k,j}]\\ +\frac{\tau L^{2}\alpha^{2}}{2}(1+\rho\tau)\frac{1}{\sqrt{v_{k-1,j}}}\left\lVert\frac{1}{\sqrt{v_{k-1}}}\odot g_{k-1}\right\rVert^{2}.$ (B.10)
Using (B.10) and summing (B.9) over $k$ from $1$ to $K$, we get
@@
by Jensen’s inequality. Then, by Lemma B.4, we have the following inequality
@@
By Lemma B.2 and taking conditional expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\alpha^{2}\frac{L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (B.11)
Let us recall from (B.4) that we have
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle=\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle.$ (B.12)
The second order term in (B.11) is equal to
@@
Also, following the proof of Theorem 5.1, the second term in (B.12) can be upper bounded by
@@
$-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}$ $\leqslant(\alpha+L_{1}\alpha^{2})\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right]$
$\leqslant 2\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right],$ (B.13)
@@
By the A.M.-G.M. inequality (i.e., $2ab\leqslant\frac{1}{\rho_{1}}a^{2}+\rho_{1}b^{2}$ for any $\rho_{1}\in\mathbb{R}_{++}$), we have
@@
$\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\right]$ $\leqslant\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]$
@@
$\left\lVert\nabla F(x_{k-1})\right\rVert\leqslant\left\lVert\nabla F(x_{k-1})-\nabla F(x_{k})\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert x_{k}-x_{k-1}\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert.$
@@
$\leqslant\frac{1}{\sqrt{v_{k-1}}}\left[2(L_{0}^{2}+L_{1}^{2}\left\lVert\nabla F(x_{k})\right\rVert^{2})\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\right]+\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]$

### v2

@@ C.2.3 Technical Lemmas
Lemma C.1.
Let $(a_{k})_{k\in\mathbb{N}}\subset\mathbb{R}_{+}$ be a sequence of nonnegative real numbers with $a_{0}>0$. Then, the following inequalities hold:
@@
We state without proof a similar descent lemma for $(L_{0},L_{1})$-smooth functions. Its proof can be found in (Zhang et al., 2020a), Lemma A.3.
Lemma C.2 (Descent lemma for $(L_{0},L_{1})$-smooth functions).
@@
C.3 Convergence Results for AdaGrad-Norm and AdaGrad
C.3.1 Full Statement of Theorem 5.1
Theorem C.1 (AdAdaGrad-Norm).
Suppose that Assumption 1 holds. Let $(x_{k})_{k\in\mathbb{N}^{*}}$ be the AdaGrad-Norm iterates (10) with any step size $\alpha>0$, where the batch sizes $(b_{k})_{k\in\mathbb{N}^{*}}$ are chosen such that either the (exact variance) norm test (4) with constant $\eta\in(0,1)$ or the (exact variance) augmented inner product test (5) and (7) with constants $(\vartheta,\nu)\in\mathbb{R}_{++}^{2}$ is satisfied at each iteration. Then, with probability at least $1-\delta$, we have
The constants $c_{1}$, $c_{2}$ and $c_{3}$ are defined through $c_{1}\coloneqq\frac{2}{\alpha(1-\rho^{-1})}(F(x_{1})-F^{\star}+\tau\alpha\left\lVert\nabla F(x_{0})\right\rVert^{2}/(2\sqrt{v_{0}})+\tau L^{2}\alpha^{3}(1+\rho\tau)/\sqrt{v_{0}}-L\alpha^{2}\log(v_{0})/2)$, $c_{2}\coloneqq 2L\alpha/(1-\rho^{-1})$, and $c_{3}\coloneqq 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)$, where $\rho\in\left(1,\infty\right)$ is a constant, with $\tau=1+\eta^{2}$ for the norm test and $\tau=1+\vartheta^{2}+\nu^{2}$ for the augmented inner product test.
C.3.2 Proof of Theorem 5.1
@@ C.3 Convergence Results for AdaGrad-Norm and AdaGrad
$=\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle,$ (C.4)
@@
$\mathbb{E}_{k}[F(x_{k+1})]\\ \leqslant F(x_{k})-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (C.5)
@@
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)g_{k}\right]\right\rangle\\ \leqslant\frac{1}{2}\left(1+\frac{1}{\rho}\right)\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{\tau L^{2}\alpha^{2}}{2}(1+\rho\tau)\frac{\left\lVert g_{k}\right\rVert^{2}}{v_{k+1}^{\nicefrac{{3}}{{2}}}}.$ (C.6)
Proof of Lemma C.1.
@@
$\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}=\frac{v_{k}-v_{k-1}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}=\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k-1}}\sqrt{v_{k}}(\sqrt{v_{k}}+\sqrt{v_{k-1}})}.$ (C.7)
@@
$\leqslant\mathbb{E}_{k}\left[\frac{\left\lVert\nabla F(x_{k-1})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k}}}\right]+\frac{1}{\sqrt{v_{k-1}}}\left(L^{2}\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2L\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\right)$
@@
By the A.M.-G.M. inequality with $\rho>0$, we have
@@
$\leqslant\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{2\sqrt{v_{k-1}}}+\frac{\tau}{2}\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]+\frac{\rho\tau^{2}L^{2}\alpha^{2}}{2}\cdot\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}+\frac{1}{2\rho}\cdot\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\frac{\tau L^{2}\alpha^{2}}{2}\cdot\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}$
@@
We now continue with the proof of Theorem 5.1. Using Lemma C.1, (C.5) is upper bounded by
@@
$\mathbb{E}[F(x_{k+1})]\leqslant\mathbb{E}[F(x_{k})]-\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}-\frac{\alpha}{2}\left(1-\frac{1}{\rho}\right)\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]+\frac{\tau\alpha}{2}\mathbb{E}[\varphi_{k-1}-\varphi_{k}]\\ +\frac{L\alpha^{2}}{2}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right].$ (C.8)
Summing the last two terms on the right hand side over $k$ from $1$ to $K$ and using Lemma C.1 yields
$\frac{L\alpha^{2}}{2}\sum_{k=1}^{K}\mathbb{E}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}+\frac{\tau L^{2}\alpha^{3}}{2}(1+\rho\tau)\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert g_{k-1}\right\rVert^{2}}{v_{k-1}^{\nicefrac{{3}}{{2}}}}\right]\\ \leqslant\frac{L\alpha^{2}}{2}(\mathbb{E}\log v_{K}-\log v_{0})+\tau L^{2}\alpha^{3}(1+\rho\tau)\frac{1}{\sqrt{v_{0}}}.$
Thus, summing (C.8) over $k$ from $1$ to $K$ and rearranging yields
@@
Lemma C.2.
For any positive constants $(a,b)\in\mathbb{R}_{++}^{2}$, if $x\leqslant a+b\log x$ for any $x>0$, then $x\leqslant 2a-2b+4b\log(b/2+1)\leqslant 2a+4b\log(b/2+1)$.
Proof of Lemma C.2.
Let $h(x)\coloneqq x/2-\log x$. Then $h^{\prime}(x)=1/2-b/x=0$ if and only if $x=2b$. Since $h^{\prime\prime}(x)=b/x^{2}>0$ for any $x>0$, so $h$ achieves its minimum at $2b$, which means $h(x)\geqslant h(2b)=b-b\log(2b)$. By the A.M.-G.M. inequality, we have $b+2\geqslant 2\sqrt{2b}$ and taking logarithm yields $\log(b+2)\geqslant\log 2+\frac{1}{2}(\log 2+\log b)$. Therefore, we obtain $b\log(2b)=b(\log 2+\log b)\leqslant 2b(\log(b+2)-\log 2)=2b\log\left(b/2+1\right)$. Consequently, we get $x-b\log x=x/2+x/2-b\log x\geqslant x/2+b-b\log(2b)\geqslant x/2+b-2b\log\left(b/2+1\right)$. Therefore, we have $x/2+b-2b\log\left(b/2+1\right)\leqslant a$, which implies the desired inequality. ∎
By Lemma C.2, we have $\mathbb{E}[\sqrt{v_{k}}]\leqslant 2c_{1}\tau+8c_{2}\tau\log(c_{2}\tau+1)\eqqcolon c_{3}$.
@@
$c_{1}+2c_{2}\log\mathbb{E}[\sqrt{v_{k}}]\geqslant\sum_{k=1}^{K}\mathbb{E}\left[\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\right]\\ \geqslant\mathbb{E}\left[\frac{1}{\sqrt{v_{K}}}\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}\right]\geqslant\frac{\mathbb{E}\left[\sqrt{\sum_{k=1}^{K}\left\lVert\nabla F(x_{k})\right\rVert^{2}}\right]^{2}}{\mathbb{E}[\sqrt{v_{K}}]},$
@@
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\frac{1}{\sqrt{v_{k-1}}}\odot\nabla F(x_{k})\right\rangle\\ +\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right)\odot g_{k}\right]\right\rangle+\frac{L\alpha^{2}}{2}\mathbb{E}_{k}\left\lVert\frac{1}{\sqrt{v_{k}}}\odot g_{k}\right\rVert^{2}.$ (C.9)
@@
$\mathbb{E}_{k}\left[\partial_{j}F(x_{k})\left(\frac{1}{\sqrt{v_{k-1,j}}}-\frac{1}{\sqrt{v_{k,j}}}\right)g_{k,j}\right]\leqslant\frac{1}{2}\left(1+\frac{1}{\rho}\right)\frac{(\partial_{j}F(x_{k}))^{2}}{\sqrt{v_{k-1,j}}}+\frac{\tau}{2}\mathbb{E}_{k}[\widetilde{\varphi}_{k-1,j}-\widetilde{\varphi}_{k,j}]\\ +\frac{\tau L^{2}\alpha^{2}}{2}(1+\rho\tau)\frac{1}{\sqrt{v_{k-1,j}}}\left\lVert\frac{1}{\sqrt{v_{k-1}}}\odot g_{k-1}\right\rVert^{2}.$ (C.10)
Using (C.10) and summing (C.9) over $k$ from $1$ to $K$, we get
@@
by Jensen’s inequality. Then, by Lemma C.2, we have the following inequality
@@
By Lemma C.2 and taking conditional expectation with respect to $\mathcal{F}_{k}$, we have
$\mathbb{E}_{k}[F(x_{k+1})]\leqslant F(x_{k})-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\alpha^{2}\frac{L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}.$ (C.11)
Let us recall from (C.4) that we have
$\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle=\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}+\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\left(\frac{1}{\sqrt{v_{k}}}-\frac{1}{\sqrt{v_{k-1}}}\right)g_{k}\right]\right\rangle.$ (C.12)
The second order term in (C.11) is equal to
@@
Also, following the proof of Theorem 5.1, the second term in (C.12) can be upper bounded by
@@
$\quad\;-\alpha\left\langle\nabla F(x_{k}),\mathbb{E}_{k}\left[\frac{g_{k}}{\sqrt{v_{k}}}\right]\right\rangle+\frac{L_{1}\alpha^{2}\left\lVert\nabla F(x_{k})\right\rVert}{2}\cdot\mathbb{E}_{k}\left\lVert\frac{g_{k}}{\sqrt{v_{k}}}\right\rVert^{2}$ (C.13)
$\leqslant(\alpha+L_{1}\alpha^{2})\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right]$
$\leqslant 2\alpha\frac{\left\lVert\nabla F(x_{k})\right\rVert}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{\sqrt{v_{k}}+\sqrt{v_{k-1}}}\right],$ (C.14)
@@
By the A.M.-G.M. inequality (i.e., $2ab\leqslant\frac{1}{\rho_{1}}a^{2}+\rho_{1}b^{2}$ for any $\rho_{1}>0$), we have
@@
$\quad\;\frac{\left\lVert\nabla F(x_{k})\right\rVert^{2}}{\sqrt{v_{k-1}}}\cdot\mathbb{E}_{k}\left[\frac{\left\lVert g_{k}\right\rVert^{2}}{(\sqrt{v_{k}}+\sqrt{v_{k-1}})^{2}}\right]$
$\leqslant\left\lVert\nabla F(x_{k})\right\rVert^{2}\cdot\mathbb{E}_{k}\left[\frac{1}{\sqrt{v_{k-1}}}-\frac{1}{\sqrt{v_{k}}}\right]$
@@
$\left\lVert\nabla F(x_{k-1})\right\rVert$ $\leqslant\left\lVert\nabla F(x_{k-1})-\nabla F(x_{k})\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert$
$\leqslant(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert x_{k}-x_{k-1}\right\rVert+\left\lVert\nabla F(x_{k})\right\rVert.$
@@
$\quad\;\left\lVert\nabla F(x_{k})\right\rVert^{2}-\left\lVert\nabla F(x_{k-1})\right\rVert^{2}$
$=\left(\left\lVert\nabla F(x_{k})\right\rVert-\left\lVert\nabla F(x_{k-1})\right\rVert\right)\left(\left\lVert\nabla F(x_{k})\right\rVert+\left\lVert\nabla F(x_{k-1})\right\rVert\right)$
@@
$\leqslant\frac{1}{\sqrt{v_{k-1}}}\left[2(L_{0}^{2}+L_{1}^{2}\left\lVert\nabla F(x_{k})\right\rVert^{2})\left\lVert x_{k}-x_{k-1}\right\rVert^{2}+2(L_{0}+L_{1}\left\lVert\nabla F(x_{k})\right\rVert)\left\lVert\nabla F(x_{k})\right\rVert\left\lVert x_{k}-x_{k-1}\right\rVert\right]$
$\qquad\qquad+\mathbb{E}_{k}[\varphi_{k-1}-\varphi_{k}]$